OpenStudy (anonymous):

Have not the foggiest idea on how to complete this :( A rubber ball dropped on a hard surface takes a sequence of bounces, each one 3/5 as high as the preceding one. If this ball is dropped from a height of 10 feet, what is the total vertical distance it has traveled after it hits the surface the fifth time? Be sure to account for both drop and rebound distances.

OpenStudy (anonymous):

well, thought i had it with this tl;dr 10 + ((10*(3/5))^1*2) + ((10*(3/5))^2*2) + ((10*(3/5))^3*2) + (10*(3/5))^4 but it isn't working, doesn't match my answer choices

OpenStudy (anonymous):

All I've got are these choices and the answer to what I came up with is 1,822 which is obviously wrong.

OpenStudy (mertsj):



OpenStudy (jdoe0001):

if the ball bounces the next time 3/5 of the previous height let's say the previous height is "a" what would be the height on the next bounce?

OpenStudy (anonymous):

@jdoe0001 6

OpenStudy (jdoe0001):

well, yes... notice it's just a geometric sequence, that has a 1st term of "10" and we dunno the 5th term but we know the "common ratio"

OpenStudy (anonymous):

Is the common ratio 0.6?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

Now what do I do with that? :|

OpenStudy (jdoe0001):

well, use the geometric sequence formula, which you surely have =)

OpenStudy (anonymous):

I may have to look that up, my homework doesn't give it to me.

OpenStudy (jdoe0001):

\(\large \bf a_{\color{red}{ n}}=a_1\cdot r^{{\color{red}{ n}}-1}\) r= common ratio

OpenStudy (anonymous):

a_n=10*0.6^5-1=-0.2224? That can't be right :c

OpenStudy (anonymous):

AUGH I pulled a ditz move there, forgot the brackets *^-^ 10*0.6^(5-1)=1.296

OpenStudy (jdoe0001):

ahemm... ? so we know the common ratio say 3/5 so \(\large \bf a_{\color{red}{ n}}=a_1\cdot r^{{\color{red}{ n}}-1}\implies a_{\color{red}{ 5}}=10\cdot \left(\cfrac{3}{5}\right)^{{\color{red}{ 5}}-1}\)

OpenStudy (anonymous):

And I'm still wrong because that isn't a choice either.

OpenStudy (jdoe0001):

If this ball is dropped from a height of 10 feet,\(\bf \text{ what is the total vertical distance it has traveled}\) after it hits the surface the fifth time?

OpenStudy (jdoe0001):

you're being asked for the sequence SUM, just like the in the previous exercise

OpenStudy (anonymous):

oh, i pulled another ditz move on that one. i need to do what i tried to do the first time

OpenStudy (anonymous):

give me a bit to try and do that

OpenStudy (mertsj):

\[10+2(\frac{3}{5})(10)+2(\frac{3}{5})^2(10)+2(\frac{3}{5})^3(10)+2(\frac{3}{5})^4(10)\]

OpenStudy (anonymous):

Is it B?

OpenStudy (anonymous):

36 14/125

OpenStudy (anonymous):

And yeah the equation you came up with was what I was doing pretty much.

OpenStudy (jdoe0001):

well... what did you get?

OpenStudy (jdoe0001):

notice Mertsj 's addition

OpenStudy (anonymous):

10 + 2 (3/5) 10 + 2 (3/5)^2 10 + 2 (3/5)^3 10 + 2 (3/5)^4 10 = 36.112 = 36 14/125

OpenStudy (mertsj):

\[10+12+7\frac{1}{5}+4\frac{8}{25}+2\frac{74}{125}=36\frac{14}{125}\]

OpenStudy (anonymous):

Oh nvm, I see that it is correct. Thanks mates ;) I dunno how I could repay you guys but I would pay if I could get help like this more often.